OpenStudy (anonymous):

Find the common factor of all the terms of the polynomial 3x + 6.

OpenStudy (ash2326):

We have \[3x+6\] Let's write the factors of each term \[3\times x+ 2\times 3\] We notice both the terms have a 3, we can take it out \[3 \times(x+2)\] or \[3(x+2)\] Here the common factor is 3:)

OpenStudy (ash2326):

@shearainh2o Did you understand?